# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v4

@@ 7 Policy
In the realm of manipulation, policies can generate signals for robots’ execution based on input. The types of outputs from policies encompass code, target pose, and delta pose. Policies generate code for robot execution directly, aiding detailed observation by humans for debugging. Generating delta poses directly through policies allows for conversion into trajectory through time sequences, offering a more end-to-end approach compared to target pose, which often requires subsequent motion planning.
The policy model for outputting delta pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. Instead, it takes observation images as input and directly outputs the direction and magnitude of the next movement. While this approach is more end-to-end, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layers. RT-2 (Brohan et al. (2023)) refers to this approach as Vision-Language-Action (VLA). Following this naming convention, we divide the policy into Vision-Language-Action-Code (VLAC), Vision-Language-Action-Target-Pose (VLATP), Vision-Language-Action-Delta-Pose (VLADP) based on different output action.
@@ 7.1 VLAC
Code generation and program synthesis have been demonstrated to be capable of developing generalizable, interpretable policy (Trivedi et al. (2021). However, a robot capable of generating code for multiple tasks, rich knowledge across various domains is essential (Ellis et al. (2023)). Therefore, scholars aim to apply the prior knowledge of LLM to code generation task (Chen et al. (2021b); Austin et al. (2021)). Code-As-Policy (Liang et al. (2023)) demonstrates the possibility of using LLMs to directly generate code for robot execution based on prompts. The study shows that 1) code-writing LLMs enable novel reasoning capability, such as encoding spatial relationships by leveraging familiarity with third-party libraries and 2) hierarchical code-writing inspired by recursive summarization improves code generation. In PROGPROMPT (Singh et al. (2023)), assertions are added to the generated code. When executing assertions, environment state feedback is obtained to check if the environment satisfies the pre- and post-conditions of the task.
7.2 VLATP
@@
The utilization of foundation models to generate target pose can be categorized into three approaches: 1) Directly using existing foundation models to output target pose. 2) Training RFMs to output target pose through reinforcement learning. 3) Training RFMs to generate target pose through imitation learning.
@@
Utilizing foundation models trained on existing large-scale internet datasets enables the direct perception of observation images and outputting target poses. Instruct2Act (Huang et al. (2023b)) utilizes CLIP and SAM to identify manipulated objects within observation images and outputs the 3D position of these manipulated objects from 2D image. DALL-E-Bot (Kapelyukh et al. (2023)) employs DALL-E to generate target images for tasks and generates target poses for manipulation by combining the target image with the observation image. ReKep (Huang et al. (2024a)) uses VFM and VLM to extract relational keypoint constraints from language instructions and RGBD observations. It then applies an optimization solver to generate a series of end-effector poses.
Training via the collection of a large manipulation dataset primarily involves utilizing off-line reinforcement learning. PI-QT-Opt (Lee et al. (2023)) leverages a large-scale, multi-task dataset and employs a model-free off-policy reinforcement learning approach for training. RL@Scale (Herzog et al. (2023a)) provides extensive empirical validation through training on real-world data collected over 24 months from experimentation across a fleet of 23 robots in three office buildings in the waste sorting application. Q-Transformer (Chebotar et al. (2023a)) facilitates training high-capacity sequential architectures on mixed-quality data by applying transformer models to RL.
@@
As for imitation learning methods, CLIPort (Shridhar et al. (2021)) demonstrates the capability of imitation learning in language-conditioned general manipulation. However, CLIPort (Shridhar et al. (2021)) addresses 4-DoF end-effector pose prediction by treating it as a pixel classification problem. Keypoint-based approaches are extended to handle 6-DoF end-effector pose prediction. Due to keypoint-based methods primarily focus on 3D scene-to-action tasks, these methods become computationally expensive as resolution requirements increase (Ke et al. (2024)). To address high spatial resolution, PerAct (Shridhar et al. (2023)) uses the latent set self-attention of Perceiver (Jaegle et al. (2021)), which has linear complexity with voxels. Act3D (Gervet et al. (2023)) represents scenes as a continuous 3D feature field, transforming 2D model features into 3D feature clouds using sensed depth. It learns 3D features through recurrent coarse-to-fine point sampling.
@@
Some research has extended the work on PerAct (Shridhar et al. (2023)) and Act3D (Gervet et al. (2023)). ChainedDiffuser (Xian et al. (2023)) builds upon Act3D (Gervet et al. (2023)) by replacing the motion planner with a diffusion model. This approach addresses the challenges of continuous interaction tasks. The 3D Diffuser Actor (Ke et al. (2024)), similar to Act3D (Gervet et al. (2023)), employs tokenized 3D scene representations. However, unlike Act3D (Gervet et al. (2023)) and 3D Diffusion Policy (Ze et al. (2024b)) with 1D point cloud embeddings, 3D Diffuser Actor (Ke et al. (2024)) leverages CLIP to extract features from 2D images and aggregates them into a 3D scene representation. GNFactor (Ze et al. (2023)) improves upon PerAct (Shridhar et al. (2023)) by enhancing 3D semantic features. It achieves this by distilling pre-trained semantic features from 2D foundation models into Neural Radiance Fields (NeRFs). DNAct (Yan et al. (2024)) builds on GNFactor (Ze et al. (2023)) and transforms the perceiver model into a diffusion head. VoxAct-B (Liu et al. (2024b)) uses VLM to divide the task into subtasks for the left arm and the right arm and applies PerAct (Shridhar et al. (2023)) to generate separate target poses for each arm.
@@
RFMs based on the above work are also under development. LEO (Huang et al. (2023a)) expands upon language foundation models by incorporating modalities like images and 3D point clouds. It fine-tunes manipulation datasets using the LoRA method. This showcases the ability to transfer the original foundation model to more modalities and manipulation tasks. VIMA (Jiang et al. (2023)) and MIDAS (Li et al. (2023c)) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs. Xu et al. (2024) considers the task goal, the object’s physical properties, and the end-effector’s design and creates a ManiFoundation model to generate the target pose. However, the target pose output by the ManiFoundation model is not 6D pose. Instead, it provides the positions of multiple contact points and the force to be applied at each contact point.
@@
Figure 7: Foundation Models for Policy. VLAC integrates task instruction, pre-written APIs, and example inputs into LLM. It generates corresponding execution code. When the code is executed, feedback from the environment is obtained. This feedback checks if the environment meets the task’s pre- and post-conditions (Singh et al. (2023)). 2) VLATP without foundation models fine-turning inputs task instruction into LLM, which specifies the manipulated object. The observation image is fed into VFM for object segmentation, and both the manipulated object and object segmentation images are input into VLM, outputting the pixel mapping of the object to be manipulated into cartesian 3D target pose (Huang et al. (2023b)). 3) VLATP with foundation models fine-tuning inputs task instruction and multimodal perception into a pre-trained model such as LLM or VLM after fine-tuning to output target pose (Huang et al. (2023a)). 4) VLADP without fine-turning. LLM outputs the description of the object to be manipulated based on task instruction. VLM then outputs the BBox of the corresponding object in RGB. Next, segmentation is performed using VFM. 3D Value Map is updated based on depth information, generating delta pose (Huang et al. (2023c)). 5) VLADP outputs delta pose and the predicted observation image after action execution by inputting task instruction and multimodal perception into a pre-trained model after fine-tuning. The difference between VLADP and VLATP with fine-tuning lies in generating delta poses directly through policies, allowing for conversion into trajectory through time sequences, offering a more end-to-end approach compared to target pose. Target pose often requires subsequent motion planning. Outputting delta pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. However, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layer. As for imagining the image after the next movement, predicting both the next action and the hindsight image can improve the performance (Bousmalis et al. (2023)). 6)Foundation Models assisting for Reinforment Learning. LLM generates subgoals based on task instruction to transform long-horizon tasks into short-horizon ones (Di Palo et al. (2023)), facilitating RL learning. LLM also creates a reward function for RL according to task instruction (Ma et al. (2023b)), while VLM can utilize prior knowledge to provide predicted action and sparse/dense reward, enhancing the effective exploration in reinforcement learning.(Ye et al. (2023a)).
@@ 7.3 VLADP
RoboNet (Dasari et al. (2019)) establishes an initial pool of 15 million video frames from 7 different robot platforms, aiming to learn a generalizable model for vision-based robotic manipulation through an end-to-end approach. This learning method closely mimics human learning, eliminating the need for calibration and 2D-to-3D mapping to understand spatial 3D coordinates. Due to imagination, RoboNet generates predicted 2D images of the next movement and estimates which pixel the robot should move to in the next step, ultimately converting it into the action the robot should take. This concept of imagination is applied in both 3D-VLA (Zhen et al. (2024)) and SuSIE (Black et al. (2023)).
RobotNet can only achieve one-model corresponding to one-task. One-model corresponding to one-task methods have also achieved significant progress. ACT (Zhao et al. (2023a)) uses action chunks to reduce compounding errors. Diffusion policy (Chi et al. (2023)) applies the idea of diffusion to visuo-motor control, tackling challenges such as action multimodality and sequential correlation to handle high-dimensional action sequences.
@@
Effective robotic multi-task learning necessitates a high-capacity model, hence Gato (Reed et al. (2022)) and RT-1 (Brohan et al. (2022)) devise transformer-based architectures. Nonetheless, RT-1 and Gato differ; RT-1’s input lacks proprioception from the robot body, while Gato incorporates proprioception. Though there’s no evidence indicating the superiority of either input method, intuitively, solely observing through images may compromise stability by omitting proprioception. Building upon Gato, RoboCat (Bousmalis et al. (2023)) demonstrates that a large sequence model can learn unseen tasks through few-shot learning. It proposes a simple but effective self-improvement process. Additionally, it shows that predicting both the next action and the hindsight image after executing that action can enhance performance. Building upon RT-1, RoboAgent (Bharadhwaj et al. (2023)) enhances model generalization and stability through data augmentation and action-chunking. MOO (Stone et al. (2023)) leverages Owl-ViT to extract object locations from observation images, enhancing RT-1’s open-set detection capability.
@@
Utilizing pre-trained VLMs (Zhang et al. (2024b)) for fine-tuning to construct RFMs is considered efficient. RT-2 (Brohan et al. (2023)) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (Chen et al. (2023b)) and PaLM-E (Driess et al. (2023)) after treating delta pose as tokens. However, this approach necessitates a lot of data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (Padalkar et al. (2023b)) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). However, the current Open X-Embodiment dataset faces the heterogeneity dataset challenge. Octo (Team et al. (2024)), RDT-1B (Liu et al. (2024d)), HPT (Wang et al. (2024c)) and PI (Black et al. (2024)) propose multi-module networks to address this issue. RT-H (Belkhale et al. (2024)) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses and human intervention can enable robots to adjust trajectories based on human language instruction. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (Li et al. (2023d)) and OpenVLA (Kim et al. (2024)) emerge, effectively enabling a robot manipulation policy with VLMs.
@@
Internet videos contain information on the physics and dynamics of the world, some studies have explored training VLMs using both video datasets and manipulation data. GR-1 (Wu et al. (2023a)) is initially trained on a large-scale video dataset for video prediction, and then seamlessly fine-tuned with manipulation data. GR-2 (Cheang et al. (2024)) uses VQGAN to convert each image into discrete tokens and is trained with a larger text-video dataset than GR-1. LAPA (Ye et al. (2024)) begins by extracting the latent delta action between video frames. It then labels the video dataset with this information. These labeled datasets are used to train a VLM network. Finally, a small-scale robot manipulation dataset is applied for fine-tuning, enabling the mapping of latent delta actions to robot actions. At the same time, BC-Z (Jang et al. (2022)) and Vid2Robot (Jain et al. (2024)) introduce a video-conditioned policy that uses video as task instructions. MUTEX (Shah et al. (2023)) extends instruction to various modalities and develops speech-conditioned, speech-goal-conditioned, image-goal-conditioned, and text-goal-conditioned.
Previous studies, such as GR-1 (Wu et al. (2023a)) and GR-2 (Cheang et al. (2024)), train the policy head using MSE regression. In contrast, OpenVLA (Kim et al. (2024)) and RT-2 (Brohan et al. (2023)) apply classification loss for their policy head. Building on the success of diffusion policy (Chi et al. (2023)), TinyVLA (Wen et al. (2024b)) adopts a diffusion head as its policy head, achieving better performance than OpenVLA.
Utilizing pre-trained foundation models without fine-tuning reduces the data collection and training steps. Voxposer (Huang et al. (2023c)) utilizes LLMs generate code that interacts with VLMs to produce value maps affordance maps and constraint maps, collectively referred to as value maps, grounded in the robot’s observation space. These composed value maps serve as objective functions for motion planners to synthesize trajectories for robot manipulation.

### v5

@@ 7 Policy
The policy is divided into two categories: object/action-centric methods and end-to-end methods. Object/action-centric methods extract attributes from observations, such as bounding boxes, masks (Sajjan et al. (2020)), or 3D spatial action-value map (Shi et al. (2024)). These extracted attributes are then transformed into either a sequence of key poses or a single key pose, which is used in motion planning to guide robot motion. End-to-end methods directly map observation to robot action (Chi et al. (2023)). They eliminate the need for attribute extraction.
End-to-end methods are mainly divided into reinforcement learning (Herzog et al. (2023)) and imitation learning (Dasari et al. (2019)). Recent end-to-end methods have made significant progress. ACT (Zhao et al. (2023a)) uses action chunks to reduce compounding errors. Diffusion policy (Chi et al. (2023)) applies the idea of diffusion to visuo-motor control, tackling challenges such as action multimodality and sequential correlation to handle high-dimensional action sequences.
However, the above methods are all one-model-for-one-task, lacking general-purpose capability. Due to the development of foundation models, general-purpose models have advanced. The representation of task instruction can be categorized into four types: language, human video, goal image, and multimodal prompts. BC-Z (Jang et al. (2022)) and Vid2Robot (Jain et al. (2024)) introduce a video-conditioned policy that uses human video as task instructions. DALL-E-Bot (Kapelyukh et al. (2023)) employs DALL-E to generate target images for tasks and generates actions for manipulation by combining the target image with the observation image. VIMA (Jiang et al. (2023)) and MIDAS (Li et al. (2023c)) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs. MUTEX (Shah et al. (2023)) extends instruction to various modalities and develops speech-conditioned, speech-goal-conditioned, image-goal-conditioned, and text-goal-conditioned.
Language-conditioned general-purpose policies remain the predominant paradigm in current research. RT-2 (Brohan et al. (2023)) refers to this approach as Vision-Language-Action (VLA). Following this naming convention, we divide the policy into Vision-Language-Action-Key-Pose (VLAKP), Vision-Language-Action-Dense-Pose (VLADP). VLAKP is more similar to traditional object/action-centric approaches. On the other hand, VLADP resembles traditional end-to-end methods. Recent studies have explored the use of foundation models to directly synthesize low-level policy code (Liang et al. (2023); Yoshida et al. (2025)). Such policies generate executable code for robots, enabling fine-grained human inspection and debugging. We denote this paradigm as Vision-Language-Action-Code (VLAC).
The above methods integrate foundation models into policy models to guide action generation. Currently, some approaches leverage foundation models to assist in reinforcement learning training.
@@ 7.1 VLAC
Code generation and program synthesis have been demonstrated to be capable of developing generalizable, interpretable policy (Trivedi et al. (2021). However, a robot capable of generating code for multiple tasks, rich knowledge across various domains is essential (Ellis et al. (2023)). Therefore, scholars aim to apply the prior knowledge of LLM to code generation task (Chen et al. (2021b); Austin et al. (2021)). Code-As-Policy (Liang et al. (2023)) demonstrates the possibility of using LLMs to directly generate code for robot execution based on prompts. The study shows that 1) code-writing LLMs enable novel reasoning capability, such as encoding spatial relationships by leveraging familiarity with third-party libraries and 2) hierarchical code-writing inspired by recursive summarization improves code generation. From-text-to-motion (Yoshida et al. (2025)) translates descriptions of human actions into humanoid robot motion code, enabling it to perform various tasks autonomously.
7.2 VLAKP
@@
The utilization of foundation models to generate key poses can be categorized into two approaches: 1) Directly using existing foundation models to output key poses. 2) Training RFMs to generate key poses through imitation learning.
@@
Utilizing foundation models pre-trained on existing large-scale internet datasets enables the direct perception of observation images and outputting key poses. Instruct2Act (Huang et al. (2023b)) utilizes CLIP and SAM to identify manipulated objects within observation images and outputs the 3D position of these manipulated objects from 2D image. Voxposer (Huang et al. (2023c)) utilizes LLMs generate code that interacts with VLMs to produce affordance maps and constraint maps, collectively referred to as value maps, grounded in the robot’s observation space. These composed value maps serve as objective functions for motion planners to synthesize trajectories for robot manipulation. ReKep (Huang et al. (2024a)) uses VFM and VLM to extract relational keypoint constraints from language instructions and RGBD observations. It then applies an optimization solver to generate a series of end-effector poses.
@@
As for imitation learning methods, CLIPort (Shridhar et al. (2021)) demonstrates the capability of imitation learning in language-conditioned general manipulation. However, CLIPort (Shridhar et al. (2021)) addresses 4-DoF end-effector pose prediction by treating it as a pixel classification problem. Keypoint-based approaches are extended to handle 6-DoF end-effector pose prediction. Due to keypoint-based methods primarily focus on 3D scene-to-action tasks, these methods become computationally expensive as resolution requirements increase (Ke et al. (2024)). To address high spatial resolution, PerAct (Shridhar et al. (2023)) uses the latent set self-attention of Perceiver (Jaegle et al. (2021)), which has linear complexity with voxels. Act3D (Gervet et al. (2023)) represents scenes as a continuous 3D feature field, transforming 2D model features into 3D feature clouds using sensed depth and learns a 3D feature field of arbitrary spatial resolution through recurrent coarse-to-fine point sampling.
@@
Some research has extended the work on PerAct (Shridhar et al. (2023)) and Act3D (Gervet et al. (2023)). ChainedDiffuser (Xian et al. (2023)) builds upon Act3D (Gervet et al. (2023)) by replacing the motion planner with a diffusion model. This approach addresses the challenges of continuous interaction tasks. The 3D Diffuser Actor (Ke et al. (2024)), similar to Act3D (Gervet et al. (2023)), employs tokenized 3D scene representations. However, unlike Act3D (Gervet et al. (2023)) and 3D Diffusion Policy (Ze et al. (2024b)) with 1D point cloud embeddings, 3D Diffuser Actor (Ke et al. (2024)) leverages CLIP to extract features from 2D images and aggregates them into a 3D scene representation. GNFactor (Ze et al. (2023)) improves upon PerAct (Shridhar et al. (2023)) by enhancing 3D semantic features. It achieves this by distilling pre-trained semantic features from 2D foundation models into Neural Radiance Fields (NeRFs). DNAct (Yan et al. (2024)) builds on GNFactor (Ze et al. (2023)) and transforms the perceiver model into a diffusion head. VoxAct-B (Liu et al. (2024b)) uses VLM to divide the task into subtasks for the left arm and the right arm and applies PerAct (Shridhar et al. (2023)) to generate separate key poses for each arm.
@@
The current imitation learning approaches also include methods using large-scale LLM/VLM. LEO (Huang et al. (2023a)) expands upon language foundation models by incorporating modalities like images and 3D point clouds. It fine-tunes manipulation datasets using the LoRA method. This showcases the ability to transfer the original foundation model to more modalities and manipulation tasks. Xu et al. (2024) considers the object motion produced by LLM/VLM, the object’s physical properties, and the end-effector’s design and creates a ManiFoundation model to generate the key pose. However, the key pose output by the ManiFoundation model is not 6D pose. Instead, it provides the positions of multiple contact points and the force to be applied at each contact point. 3D-VLA (Zhen et al. (2024)) can generate the final state image and point cloud based on user input. This goal state is then used to create key poses in 3D VLA.
@@
Figure 7: Foundation Models for Policy. VLAC integrates task instruction, pre-written APIs, and example inputs into LLM. It generates corresponding execution code (Liang et al. (2023); Yoshida et al. (2025)). 2) VLAKP without foundation models training inputs task instruction into LLM, which specifies the manipulated object. The observation image is fed into VFM for object segmentation or keypoints, and both the manipulated object and object feature images are input into VLM, outputting the pixel mapping of the object to be manipulated into cartesian 3D key pose (Huang et al. (2023b)) or relational keypoint constraints (Huang et al. (2024a)). 3) VLAKP with training uses LLM to extract tokens from Task Instructions. VFM or VLM extract features from Multi-View Images and generate 3D features through 3D Lift. Finally, these features, along with proprioception, are used as input to the Diffusion Model to generate the key pose (Ke et al. (2024)). 4) VLADP with training (Single-System) outputs dense pose and the hindsight image by inputting task instruction and observation into a pre-trained model after training (Black et al. ). The difference between VLADP and VLAKP with training lies in generating dense poses directly through policies, allowing for conversion into trajectory through time sequences, offering a more end-to-end approach compared to key pose. Key pose often requires subsequent motion planning. Outputting dense pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. However, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layer. As for imagining after the next movement, predicting both the next action and the hindsight image can improve the performance (Bousmalis et al. (2023)). 5) The VLADP with training (Double-System) generates dense pose by using models with different inference frequencies (Figure.ai (2025)). This method can effectively leverage prior knowledge from VLM and improve inference frequency. 6)Foundation Models assisting for Reinforment Learning. LLM generates subgoals based on task instruction to transform long-horizon tasks into short-horizon ones (Di Palo et al. (2023)), facilitating RL learning. LLM also creates a reward function for RL according to task instruction (Ma et al. (2023b)), while VLM can utilize prior knowledge to provide predicted action and sparse/dense reward, enhancing the effective exploration in reinforcement learning.(Ye et al. (2023a)).
Table 1: Strengths and Limitations of VLAC, VLAKP, and VLADP
Policy Strengths Limitations
VLAC There are many Internet code datasets available. These can serve as priors for code generation. This helps improve the generalization of VLAC-style methods (Liang et al. (2023)). VLAC facilitates human attention to fine-grained details, enhancing the debugging process. VLAC is able to generate more creative actions, as shown in Yoshida et al. (2025). The use of third-party libraries in VLAC can constrain the extensibility of policy tasks. Syntax errors and similar simple faults tend to appear frequently..
VLAKP Compared to VLADP, VLAKP has a key advantage. It can leverage rich prior knowledge from foundation models in CV and NLP. This supports object-centric and action-centric action generation. As a result, VLAKP requires less training data and generalizes better than VLADP (Ke et al. (2024)). VLAKP improves upon VLAC by replacing sequential chaining with spatial composition under joint optimization. This flexibility supports more diverse manipulation tasks and ensures more stable execution (Huang et al. (2023c)). Mostly relies on 3D inputs, sensitive to pose errors. Requires precise scene reconstruction and calibration. VLAKP often requires motion planning. However, motion planning is not always reliable.
VLADP Compared with VLAKP, it does not require calibration or motion planning. It is more end-to-end, making it easier to transfer learned priors to other tasks. Compared with other methods, VLADP has a smaller data annotation burden. It relies on large-scale training data, but faces data scarcity issues. It typically has a large number of parameters, requiring significant computational resources and leading to high latency. It shows poor generalization across different embodiments.
@@ 7.3 VLADP
The policy model for outputting dense pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. Instead, it takes observation images as input and directly outputs the direction and magnitude of the next waypoint. While this approach is more end-to-end, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layers.
@@
Effective robotic multi-task learning necessitates a high-capacity model, hence Gato (Reed et al. (2022)) and RT-1 (Brohan et al. (2022)) devise transformer-based architectures. Nonetheless, RT-1 and Gato differ; RT-1’s input lacks proprioception from the robot body, while Gato incorporates proprioception. Building upon Gato, RoboCat (Bousmalis et al. (2023)) demonstrates that a large sequence model can learn unseen tasks through few-shot learning. It proposes a simple but effective self-improvement process. Additionally, it shows that predicting both the next action and the hindsight image after executing that action can enhance performance. Building upon RT-1, RoboAgent (Bharadhwaj et al. (2023)) enhances model generalization and stability through data augmentation and action-chunking. MOO (Stone et al. (2023)) leverages Owl-ViT to extract object locations from observation images, enhancing RT-1’s open-set detection capability.
As for multi-task reinforcement learning. PI-QT-Opt (Lee et al. (2023)) leverages a large-scale, multi-task dataset and employs a model-free off-policy reinforcement learning approach for training. Q-Transformer (Chebotar et al. (2023)) facilitates training high-capacity sequential architectures on mixed-quality data by applying transformer models to RL.
@@
Utilizing pre-trained VLMs (Zhang et al. (2024b)) for fine-tuning to construct RFMs is considered efficient. RT-2 (Brohan et al. (2023)) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (Chen et al. (2023b)) and PaLM-E (Driess et al. (2023)) after treating pose as tokens. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (Li et al. (2023d)) and OpenVLA (Kim et al. (2024)) emerge, effectively enabling a robot manipulation policy with VLMs.
However, this approach necessitates a lot of data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (Padalkar et al. (2023b)) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). However, the current Open X-Embodiment dataset faces the heterogeneity dataset challenge. Octo (Team et al. (2024)) and HPT (Wang et al. (2024c)) propose multi-module networks to address this issue. RDT-1B (Liu et al. (2024d)) and PI (Black et al. (2024)) propose unified action space to address this issue. The Open X-Embodiment dataset lacks in-the-wild scenes. DROID (Khazatsky et al. (2024)) introduces an “in-the-wild” robot manipulation dataset. It contains 76k trajectories, or 350 hours of interaction data. This data is collected from 564 scenes, 86 tasks, and 52 buildings over 12 months.
@@
Internet videos contain information on the physics and dynamics of the world, some studies have explored training foundation models using both video datasets and manipulation data. SuSIE (Black et al. (2023)) uses an image-editing diffusion model. This model is fine-tuned on human videos and robot rollouts. It acts as a high-level planner by proposing intermediate subgoals. These subgoals can be accomplished by a low-level controller. The two training steps of SuSIE do not share weights. GR-1 (Wu et al. (2023a)) is initially trained on a large-scale video dataset for video prediction, and then seamlessly fine-tuned with manipulation data. GR-2 (Cheang et al. (2024)) uses VQGAN to convert each image into discrete tokens and is trained with a larger text-video dataset than GR-1. LAPA (Ye et al. (2024)) begins by extracting the latent delta action between video frames. It then labels the video dataset with this information. These labeled datasets are used to train a VLM network. Finally, a small-scale robot manipulation dataset is applied for fine-tuning, enabling the mapping of latent actions to robot actions.
Previous studies, such as GR-1 (Wu et al. (2023a)) and GR-2 (Cheang et al. (2024)), train the policy head using MSE regression. In contrast, OpenVLA (Kim et al. (2024)) and RT-2 (Brohan et al. (2023)) apply next-token prediction for their policy head. Building on the success of diffusion policy (Chi et al. (2023)), PI0 (Black et al. ) and TinyVLA (Wen et al. (2024b)) adopt diffusion head as its policy head, achieving better performance than OpenVLA. In order to address higher degree of multi-modality in the distribution of feasible actions for bimanual manipulation, RDT-1B (Liu et al. (2024d)) utilizes Diffusion Transformers (DiTs) as its scalable backbone network. Go-1 (Bu et al. (2025)) trains latent actions using human videos and combines those with manipulation data. Then, it trains the VLA on the merged dataset to boost the model’s generalization.
Although the diffusion policy can represent complex continuous action distributions, OpenVLA-OFT (Kim et al. (2025)) has shown in dual-arm tasks that fine-tuning the VLA with an L1 regression objective achieves performance similar to diffusion-based fine-tuning. However, it offers faster training convergence and inference speed. FAST (Pertsch et al. (2025)) proposes a new compression-based tokenization scheme for next-token prediction. This method matches the performance of diffusion VLAs, while reducing training time by up to 5x across multiple dexterous manipulation tasks.
Previous methods face a basic tradeoff: VLM backbones are general but slow, while robot visuomotor policies are fast but not general. Synchronizing both does not improve inference speed. Helix (Figure.ai (2025)) and Groot N1 (Bjorck et al. (2025)) overcome this tradeoff with two asynchronous complementary systems, trained end-to-end to communicate. However, Groot N1 makes more use of human video latent action and simulated data compared to Helix.